Circular No. 8306

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

NOVA SAGITTARII 2004 The Central Bureau has received word of the discovery of a possible

nova independently by Hideo Nishimura (Kakegawa, Shizuoka-ken, Japan; via S. Nakano, Sumoto, Japan) and by William Liller (Viña del Mar, Chile). Nishimura reported the position of the new object (mag 9.4) as α = 18^{h}19^{m}28^{s}, δ = −28^{o}35′34′′ (equinox 2000.0) from two Fuji T-Max films taken around Mar. 15.82 UT with a Pentax 200-mm f/4 lens; nothing was visible at this location on his patrol films taken on Mar. 12.82 (limiting mag 10.5). Liller found the new star at mag ∼ 8.2 on two Technical Pan films taken with an 85-mm Nikon camera lens (+ red filter) on Mar. 17.342 and 17.346, providing the position as α = 18^{h}19^{m}.6, δ = −28^{o}35′; he adds that nothing is visible at this position on his exposures taken on Feb. 26.38 (limiting mag 11.0). Following requests by the Central Bureau and the AAVSO, D. West (Mulvane, KS) obtained a CCD image showing the apparent nova at V = 8.38 ± 0.07 on Mar. 17.489, providing the following precise position: α = 18^{h}19^{m}32^{s}.29, δ = −28^{o}36′35′′.7. West adds that nothing is visible at this location on a red Digitized Sky Survey image (limiting mag ≈ 17.5). Visual magnitude estimates, provided in part by E. Waagen, AAVSO: Mar. 17.556, 8.1 (R. Royer, Springville, CA); 17.631, 8.5 (J. Bedient, Honolulu, HI); 17.821, 8.7 (A. Pearce, Nedlands, W. Australia).

IRAS 05436–0007 D. K. Ojha, N. Kusakabe, and M. Tamura, National Astronomical Ob-

servatory, Tokyo, report on observations made at the Infrared Survey Facility 1.4-m telescope at the South African Astronomical Observatory. The ‘JHK_{s}-band simultaneous camera’ (SIRIUS) was used to obtained the following magnitudes (aperture radius ∼ 8′′) for the near-infrared counterpart to IRAS 05436–0007 (cf. IAUC 8284): Feb. 23 UT, J = 10.81, H = 8.96, K_{s} = 7.41; Feb. 29, J = 10.92, H = 9.12, K_{s} = 7.62. The cometary infrared nebula is clearly seen extending toward the north from the near-infrared source. The nebula is blue (bright in J) and has a cavity structure with two rims extending toward the northeast and the northwest. The northeast rim, which can be traced out to ∼ 30′′, is brighter and sharp compared to the diffuse northwest rim.

2004 March 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green